OpenStudy (anonymous):

how do i find t^-1 (x) if t(x) = x^2 - 6x +5

OpenStudy (jennithemeani):

First you make t(x) y then switch the variables x and y and isolate y then make y t(x) again

Parth (parthkohli):

The problem with this function is that it is not one-to-one.

OpenStudy (anonymous):

yeah i got (x-1) (x-5) = y.... i need to make x the subject of the formula... i don't know how to do that

OpenStudy (lynfran):

let t(x)=y so y=x^2-6x+5 solve for x y-5+9=x^2-6x+9 y+4=(x-3)^2 sq.rt(y+4)=(x-3) sq.rt(y+4)+3=x now to find the inverse replace x with y ..and y with x so sq.rt(x+4)+3=y

OpenStudy (anonymous):

thank u

OpenStudy (lynfran):

welcome